Wayne Enterprises Inc. pays a regular annual dividend on its common shares, which is expected to grow annually in perpetuity at the rate of 3%. The dividend was $3.00 per share. Ignoring settlement, taxes and other institutional issues, what is a fair price for the stock today if investors expect an annual return of 9%? Show how you calculated your answer.
Answer #1

Next year dividend = Current dividend (1 + g)

Next year dividend = 3 (1 + 0.03)

Next year dividend = 3.09

Fair price = D1 / required rate - growth rate

Fair price = 3.09 / 0.09 - 0.03

Fair price = 3.09 / 0.06

Fair price = $51.50